OpenStudy (anonymous):

Quadratic formula, find the solution to 2/(1-x) + 3/(1+x) =1/(x-7).

hero (hero):

Quadratic formula not necessary to solve this

hero (hero):

Would you like to see the steps?

OpenStudy (anonymous):

yes please, because the ans. is reduces to 12x-35=0, so x=3.?

hero (hero):

Yes that is correct

hero (hero):

Do you still need the steps?

hero (hero):

Actually, it reduces to 12x - 35 = 1

OpenStudy (anonymous):

I don't understand how to solve the problem because the dens., confuse me.

hero (hero):

0. Given: \(\large\frac{2}{1-x} + \frac{3}{1+x} = \frac{1}{x-7}\) 1. Multiply the first fraction by (1+x)/(1+x) to get \(\large\frac{2+ 2x}{1-x^2} + \frac{3}{1+x} = \frac{1}{x-7}\) 2. Multiply the second fraction by (1-x)/(1-x) to get: \(\large\frac{2+ 2x}{1-x^2} + \frac{3-3x}{1-x^2} = \frac{1}{x-7}\) 3. Combine fractions on the left side to get: \(\large\frac{5-x}{1-x^2} = \frac{1}{x-7}\) 4. Cross Multiply to get: \((5-x)(x-7) = 1-x^2\) 5. Multiply the left side to get: \(5x-x^2 + 7x - 35 = 1-x^2\) 6. Add x^2 to both sides to get: \(5x + 7x - 35 = 1\) 7. Combine like terms to get: \(12x - 35 = 1\) 8. Finish the rest

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

I got it, thank you